OpenStudy (anonymous):

dy/dx=(8x^(3))*e^(-2y) ; y(1)=0

sam (.sam.):

Separable \[\Large \frac{dy}{dx}=(8x^{3})e^{-2y} \\ \\ \Large \int\limits \frac{1}{e^{-2y}}dy=8\int\limits x^3dx\] \[\Large \int\limits\limits e^{2y}dy=8\int\limits\limits x^3dx\] Power rule for both integral

OpenStudy (anonymous):

Thanks